How much should be invested at 8% compounded semiannually in order to have $5000 at the end of 8 years?
a. $4215.27
b. $9364.91
c. $2669.54
d. $3048.78
: How much should be invested at 8% compounded semiannually in order to have $5000 at the end of 8 years?
a. $4215.27
b. $9364.91
c. $2669.54
d. $3048.78

Answer by vleith(967)

The future value of money using compound interest is given by:
F = P(1+(r/n))^(nt)
Where
F is the future value
P is the present value
r s the annual interest rate
n is the number of compounding per year
t is the total number of compoundings
This problem asks for the present value (P)of a future sum (F=5000), when the interest is 8%(r=0.08) compounded semi-annually (n=2) for a length of 8 years (t = 8*n = 16).
Solving for P:
F = P(1+(r/n))^(nt)
5000 = P(1+(0.08/2))^(2*8)
5000 = P(1+(0.04))^(16)
5000 = P(1.04)^(16)
5000/((1.04)^(16)) = P
5000/1.872 = P
2669.54 ~=P
The answer is C
Read this for more info -->http://www.purplemath.com/modules/expofcns4.htm